### v3

# One Brookings Drive St. Louis, MO 63130 USA

Abstract
Rational inner functions are a generalization of finite Blaschke products to several variables. In this article we survey a variety of results about rational inner functions related to interpolation, sums of squares formulas, and boundary behavior. We mostly focus on two variables however in the final section we discuss higher dimensions.
Keywords:
Rational inner functions, Blaschke products, stable polynomials, rational singularities, von Neumann inequality, multivariable operator theory, sums of squares, interpolation
0.1 Introduction
Rational inner functions are a multivariable generalization of finite Blaschke products. Blaschke products are fundamental in one complex variable especially in the case of operator related function theory on the unit disk. There has even been an entire book written about them BBook . While rational inner functions in several variables cannot possibly exhibit the centrality that finite Blaschke products possess, they represent an exciting area of study overlapping with the theory of stable polynomials, multivariable operator theory, function spaces on domains in several variables, and real algebraic geometry. We hope to illustrate this in the present survey by reviewing connections to interpolation, sums of squares formulas connected to operator theory, and the study of boundary behavior of analytic functions. A great deal is known at this point, especially in two variables, but there remains a lot to be discovered.
Recall that a finite Blaschke product is a function $B$, analytic on the unit disk $\mathbb{D}=\{z\in\mathbb{C}:|z|<1\}$ of the form
$B(z)=\mu\prod_{j=1}^{n}\frac{z-a_{j}}{1-\bar{a}_{j}z}$
where $a_{1},\dots,a_{n}\in\mathbb{D}$ and $\mu$ belongs to the unit circle $\mathbb{T}=\{z\in\mathbb{C}:|z|=1\}$. Finite Blaschke products are of course rational, analytic on $\mathbb{D}$, and they are inner, meaning
$|B(z)|=1\text{ for almost every }z\in\mathbb{T}.$
The use of “almost every” here is unnecessary because $B$ is analytic on a neighborhood of the closed unit disk. This motivates our definition of a rational inner function in several variables. Our focus will be on the domain
$\mathbb{D}^{d}=\{z=(z_{1},\dots,z_{d})\in\mathbb{C}^{d}:|z_{1}|,\dots,|z_{d}|<1\}$
called the (unit) polydisk in $\mathbb{C}^{d}$ and its distinguished boundary
$\mathbb{T}^{d}=\{z=(z_{1},\dots,z_{d})\in\mathbb{C}^{d}:|z_{1}|=\dots=|z_{d}|=1\}$
called the $d$-torus.
Definition 1
Given polynomials $p,q\in\mathbb{C}[z_{1},\dots,z_{d}]$ with no common factors, where $p(z)\neq 0$ for $z\in\mathbb{D}^{d}$, the function $f=q/p$ is a rational inner function if $|f(z)|=1$ for almost every $z\in\mathbb{T}^{d}$.
The first theorem in the subject is the following characterization of rational inner functions. It helps to define a natural reflection of a polynomial $p\in\mathbb{C}[z_{1},\dots,z_{d}]$ to be
$\tilde{p}(z)=z^{n}\overline{p(1/\bar{z}_{1},\dots,1/\bar{z}_{d})}$
where $n=\deg(p):=(\deg_{z_{1}}(p),\dots,\deg_{z_{d}}(p))$ is the multi-degree of $p$. We are using multi-index notation to define $z^{n}$.
Theorem 0.1.1
Given $p,q\in\mathbb{C}[z_{1},\dots,z_{d}]$ with no common factors, where $p(z)\neq 0$ for $z\in\mathbb{D}^{d}$, the function $f=q/p$ is a rational inner function if and only if $q(z)$ is of the form
$az^{m}\tilde{p}(z)$
where $a\in\mathbb{T}$, $m\in\mathbb{Z}_{\geq 0}^{d}$ is a $d$-tuple of non-negative integers.
The converse is a consequence of $|\tilde{p}|=|p|$ on $\mathbb{T}^{d}$. Thus, every rational inner function is of the form
$z^{m}\frac{\tilde{p}(z)}{p(z)}$
for $p$ with no zeros in $\mathbb{D}^{d}$ and no factors in common with $\tilde{p}$ (the constant $a$ can be absorbed into the polynomials). We can also simply write the rational inner function as $\frac{\tilde{p}}{p}$ by reflecting $p$ at a degree higher than its multidegree. In this way we have a simple uniform way of writing rational inner functions. We shall give essentially Pfister’s proof Pfister in the next section. With Theorem 0.1.1, it is easy to construct examples of rational inner functions (that is, if one can easily construct polynomials with no zeros on $\mathbb{D}^{d}$.)
Example 1
For convenience, when working in two variables we often write $(z,w)$ instead of $(z_{1},z_{2})$. The polynomials $p_{0}(z,w)=2-z-w$ and $p_{1}(z,w)=3-z-w$ both have no zeros in $\mathbb{D}^{2}$ and have no factors in common with their reflections
$\tilde{p}_{0}(z,w)=2zw-z-w\qquad\tilde{p}_{1}(z,w)=3zw-z-w$
and therefore both lead to rational inner functions
$f_{0}(z,w)=\frac{2zw-z-w}{2-z-w}\quad f_{1}(z,w)=\frac{3zw-z-w}{3-z-w}.$
Since $p_{1}$ has no zeros on the closed bidisk $\overline{\mathbb{D}}^{2}$, $f_{1}$ extends analytically to a neighborhood of $\overline{\mathbb{D}}^{2}$. On the other hand, $p_{0}$ has a zero at $(1,1)$ and $f_{0}$ does not even extend continuously to $\overline{\mathbb{D}}^{2}$. To see this, we first note that $f_{0}$ extends continuously to $\overline{\mathbb{D}}^{2}\setminus\{(1,1)\}$ since $02-z-w=$ 0 for $|z|,|w|\leq 1$ if and only if $z=w=1$. Now, consider for each $a\in\mathbb{R}$, the limit
$\lim_{\theta\to 0}f_{0}(e^{i\theta},e^{-i\theta+ia\theta^{2}})=\lim_{\theta\to 0}\frac{ia\theta^{2}+\theta^{2}+O(\theta^{3})}{-ia\theta^{2}+\theta^{2}+O(\theta^{3})}=\frac{1+ia}{1-ia}$
which is an arbitrary element of $\mathbb{T}$ excluding $-1$. $\diamond$
Thus, the most basic difference between one and several variables is that rational inner functions in two or more variables can have boundary singularities. The example $f_{0}$ above is often called the “favorite RIF” because it is the simplest rational inner function with a boundary singularity and it exhibits a variety of interesting phenomena.
Having introduced the concept of rational inner functions, why are we interested in them? First, polynomials with no zeros on the polydisk, often called stable polynomials, are of broad interest in a variety of areas of math and rational inner functions offer a vehicle for studying them. Second, the Schur class of the polydisk consists of the analytic functions on the polydisk bounded by one and rational inner functions approximate Schur class functions locally uniformly. This makes rational inner functions an attractive family of functions to use in tackling a variety of problems because unlike with polynomials the property of being bounded by one is assured. An prominent example of this is the fact that one can prove a sum of squares formula for rational inner functions in two variables that implies an operator inequality for bounded analytic functions in two variables called Andô’s inequality as well as Agler’s Pick interpolation theorem on the bidisk.
Once rational inner functions have proven their worth we become open to studying them for their own sake. A lot of work has been done in recent years to understand the level sets of rational inner functions and the behavior of their singularities. Some of this is motivated by operator theory, some by pure function theory, and some by the natural quest for an understanding of the behavior of the zero set of a stable polynomial near a “boundary” zero.
0.2 Basic facts
We begin by establishing some basic facts about rational inner functions. We hope it is worthwhile to the reader to see some details about the fundamentals— we will be more sketchy about more modern developments. The paper of Pfister Pfister is probably the first serious study of rational inner functions and many of the results below are from this paper as well as Rudin’s book RudinBook .
Lemma 1
If $p\in\mathbb{C}[z_{1},\dots,z_{d}]$ is nonzero, then integrating with respect to normalized Lebesgue measure we have
$\int_{\mathbb{T}^{d}}\log|p|>-\infty$
and in particular, the zero set of $p$ on $\mathbb{T}^{d}$ has measure zero.
Proof
One can prove an iterated Jensen-Poisson inequality
[…]
where $d\sigma$ is normalized Lebesgue measure, $a\in\mathbb{D}^{d}$, $P_{a}(z)=\prod_{j=1}^{d}\frac{1-|a_{j}|^{2}}{|1-\bar{a}_{j}z_{j}|^{2}}$ is the Poisson kernel. Then, since $p$ is not identically zero we can find $a\in\mathbb{D}^{d}$ such that $p(a)\neq 0$ yielding
[…]
and hence $\int_{\mathbb{T}^{d}}\log|p|>-\infty$. ∎
Proof (Proof of Theorem 0.1.1)
Given $f=q/p$ with $|q/p|=1$ almost everywhere on $\mathbb{T}^{d}$ we have $|q|=|p|$ almost everywhere on $\mathbb{T}^{d}$. Let $n=\deg p,m=\deg q$. Then,
$z^{n}q(z)\tilde{q}(z)=z^{m}p(z)\tilde{p}(z)$
almost everywhere on $\mathbb{T}^{d}$. However, two polynomials that agree almost everywhere on $\mathbb{T}^{d}$ are necessarily identical on $\mathbb{C}^{d}$ (for instance, their Fourier coefficients are the polynomials coefficients). Since $p$ and $q$ are assumed to have no common factors we must have $p(z)$ divides $z^{n}\tilde{q}(z)$. Writing $z^{n}\tilde{q}(z)=p(z)g(z)$ for some factor $g$, we have
$|z^{n}\tilde{q}(z)|=|q(z)|=|p(z)|=|p(z)||g(z)|$
on $\mathbb{T}^{d}$. Therefore, $|g(z)|=1$ almost everywhere. Also, since $p$ has multidegree $n$, the multidegree of $g$ is the same as $\tilde{q}$. Set $k=\deg\tilde{q}\leq\deg q$. We have $\tilde{g}(z)g(z)=z^{k}$ on $\mathbb{C}^{d}$. Since $g(z)$ has multidegree $k$, we see that $g(z)$ is a multiple of $z^{k}$, say $g(z)=az^{k}$. We necessarily have $|a|=1$. Therefore, $z^{n}\tilde{q}(z)=az^{k}p(z)$. Reflecting at degree $m+n$ we have $q(z)=\bar{a}z^{m-k}\tilde{p}(z)$, concluding the proof. ∎
Proposition 1
Rational inner functions are bounded by $1$ in $\mathbb{D}^{d}$.
Proof
It is enough to show $|\tilde{p}(z)|\leq|p(z)|$ in $\mathbb{D}^{d}$. If $p$ has no zeros on $\overline{\mathbb{D}}^{d}$, this is immediate by the maximum principle applied to $\tilde{p}/p$. Otherwise, we consider $10<r<$ 1 and $p_{r}(z)=p(rz)$ for which $\widetilde{p_{r}}(z)=z^{n}\overline{p(r/\bar{z})}=r^{|n|}\tilde{p}(z/r)$ and
$|r^{|n|}\tilde{p}(z/r)|\leq|p(rz)|.$
Sending $r\to 1$ we see $|\tilde{p}(z)|\leq|p(z)|$. ∎
The following theorem was proven by Carathéodory in one variable; the several variable version was probably first stated and proved by Pfister Pfister ; see also RudinBook (Theorem 5.5.1).
Proposition 2
Let $f:\mathbb{D}^{d}\to\mathbb{C}$ be analytic and $|f|\leq 1$. Then, for any $r<1$ and $\epsilon>0$, there exists a rational inner function $\phi$ such that $|f(z)-\phi(z)|<\epsilon$ for $z\in r\overline{\mathbb{D}}^{d}$.
Lemma 2
Let $f:\mathbb{D}^{d}\to\mathbb{C}$ be analytic and $|f|\leq 1$. Then, for any $r<1$ and $\epsilon>0$, there exists a polynomial $P$ such that $|f(z)-P(z)|<\epsilon$ for $z\in r\overline{\mathbb{D}}^{d}$ and $|P(z)|<1$ for $z\in\overline{\mathbb{D}}^{d}$.
Proof
For each $r<1$, write $f_{r}(z)=f(rz)$. For fixed $r<1$ there exists $s<1$ such that $|f_{r}(z)-f_{rs}(z)|<\epsilon/2$ for $z\in\overline{\mathbb{D}}^{d}$ since $f_{r}$ is uniformly continuous on $\overline{\mathbb{D}}^{d}$. Note that $\sup_{\overline{\mathbb{D}}^{d}}|f_{s}|=A<1$ by the maximum principle. Choose a Taylor polynomial $P$ of $f_{s}$ so that $|f_{s}(z)-P(z)|<\min(1-A,\epsilon/2)$ for $z\in\overline{\mathbb{D}}^{d}$. Then, $|P(z)|<1$ for $z\in\overline{\mathbb{D}}^{d}$ and
$|f_{r}(z)-P_{r}(z)|\leq|f_{r}(z)-f_{rs}(z)|+|f_{rs}(z)-P_{r}(z)|<\epsilon\text{ for }z\in\overline{\mathbb{D}}^{d}.\text{ \qed}$ ∎
Proof (Proof of Proposition 2)
Choose $P$ as in Lemma 2. It is enough to show that we can approximate $P$ locally uniformly by a rational inner function. Let $n=\deg P$ be the multidegree of $P$. Since $|P|=|\tilde{P}|$ on $\mathbb{T}^{d}$, the maximum principle implies $|\tilde{P}(z)|<1$ on $\overline{\mathbb{D}}^{d}$. Let $N\in\mathbb{Z}^{d}_{>0}$. Then, $1+z^{N}\tilde{P}(z)$ has no zeros in $\overline{\mathbb{D}}^{d}$ and multidegree $n+N$. Reflecting with respect to degree $n+N$ we see
$\phi(z)=\frac{P(z)+z^{N+n}}{1+z^{N}\tilde{P}(z)}$
is a rational inner function. Note
$\phi(z)=P(z)+z^{N}\frac{z^{n}-\tilde{P}P}{1+z^{N}\tilde{P}}$
and we can make the right-most expression arbitrarily small if one of the components of $N$ is large enough since $z\in r\overline{\mathbb{D}}^{d}$.∎
0.3 Interpolation
The classical Pick interpolation theorem gives necessary and sufficient conditions when the following interpolation problem can be solved: given $z_{1},\dots,z_{n}\in\mathbb{D}$ and $t_{1},\dots,t_{n}\in\mathbb{D}$ when does there exist analytic $f:\mathbb{D}\to\mathbb{D}$ with $|f|\leq 1$ such that $f(z_{j})=t_{j}$ for $j=1,\dots,n$? The answer is if and only if
[…]
This theorem can be addressed entirely using rational inner functions (in this case, finite Blaschke products). Indeed, it is possible to show (1) implies that we have a finite Blaschke product that interpolates, and conversely, by the approximation result, Proposition 2, we only need to verify necessity for rational inner functions.
Very briefly, one way to prove sufficiency is to factor
$\frac{1-t_{j}\bar{t}_{k}}{1-z_{j}\bar{z}_{k}}=\langle u_{j},u_{k}\rangle$
for some vectors $u_{j}$ in $\mathbb{C}^{N}$ (where $N$ is the rank of the positive semi-definite matrix). After a process called a “lurking isometry” one can find a function of the form
$\phi(z)=A+zB(I-zD)^{-1}C$
that interpolates. Here $U=\begin{pmatrix}A&B\\ C&D\end{pmatrix}$ is a block unitary matrix with $A$ a $1\times 1$ matrix and $D$ an $N\times N$ matrix. It turns out functions of this form are finite Blaschke products. The lurking isometry argument is given for instance in PickBook .
To prove necessity of (1) for finite Blaschke products, one approach uses ideas from orthogonal polynomials. Given a polynomial $p\in\mathbb{C}[z]$ with no zeros on $\overline{\mathbb{D}}$ and degree $n$, we consider the measure $d\mu=\frac{1}{|p|^{2}}d\sigma$ on $\mathbb{T}$. One can show that in $L^{2}(\mu)$, $p\perp z,\dots,z^{n}$ and $\tilde{p}\perp 1,\dots,z^{n-1}$. If $q_{0},\dots,q_{n-1}$ is an orthonormal basis for $\text{span}\{z^{j}:j=0,\dots,n-1\}$ in $L^{2}(\mu)$, then $q_{0},\dots,q_{n-1},\tilde{p}$ and $p,zq_{0},\dots,zq_{n-1}$ are both orthonormal bases for $\text{span}\{z^{j}:j=0,\dots,n\}$. So, there exists a unitary matrix $U$ such that
$U\begin{pmatrix}\tilde{p}(z)\\ q_{0}(z)\\ \vdots\\ q_{n-1}\end{pmatrix}=\begin{pmatrix}p(z)\\ zq_{0}(z)\\ \vdots\\ zq_{n-1}(z)\end{pmatrix}.$
Unitarity implies that
$\tilde{p}(z)\overline{\tilde{p}(w)}+\sum_{j=0}^{n-1}q_{j}(z)\overline{q_{j}(w)}=p(z)\overline{p(w)}+z\bar{w}\sum_{j=0}^{n-1}q_{j}(z)\overline{q_{j}(w)}$
and this rearranges into
$\frac{p(z)\overline{p(w)}-\tilde{p}(z)\overline{\tilde{p}(w)}}{1-z\bar{w}}=\sum_{j=0}^{n-1}q_{j}(z)\overline{q_{j}(w)}.$
Finally, dividing by $p(z)\overline{p(w)}$ yields
$\frac{1-\phi(z)\overline{\phi(w)}}{1-z\bar{w}}=\sum_{j=0}^{n-1}\frac{q_{j}(z)}{p(z)}\frac{\overline{q_{j}(w)}}{\overline{p(w)}}.$
In particular, we immediately get (1) for $t_{j}=\phi(z_{j})$. (Note we can get the most general finite Blaschke product by allowing $\tilde{p}$ to have a monomial factor in front.) This method at first seems ad hoc but it has a variety of other purposes as the measures $\frac{1}{|p|^{2}}d\sigma$, called Bernstein-Szegő measures, can be used to solve a different type of interpolation problem, namely the truncated trigonometric moment problem; see Landau .
Everything above can be generalized to two variables if one is careful about how one generalizes.
Theorem 0.3.1 (Agler’s Pick interpolation theorem)
Given $(z_{1},w_{1}),\dots,(z_{N},w_{N})\in\mathbb{D}^{2}$ and $t_{1},\dots,t_{N}\in\mathbb{D}$, there exists an analytic function $f:\mathbb{D}^{2}\to\mathbb{D}$ with $f(z_{j},w_{j})=t_{j}$ for $j=1,\dots,N$ if and only if there exist positive semi-definite matrices $A,B$ such that
$1-t_{j}\bar{t}_{k}=(1-z_{j}\bar{z}_{k})A_{j,k}+(1-w_{j}\bar{w}_{k})B_{j,k}.$ (2)
Just as in one variable, we can prove sufficiency by solving for a rational inner function that solves the interpolation problem and it is enough to prove necessity for rational inner functions. Addressing sufficiency, one can show that if a formula like (10) holds then the interpolation problem can be solved with a function of the form
$\phi(z,w)=A+B\Delta(z,w)(I-D\Delta(z,w))^{-1}C$ (3)
where we again have a block $(1+N)\times(1+N)$ unitary $U=\begin{pmatrix}A&B\\ C&D\end{pmatrix}$ but
$\Delta(z,w)=\begin{pmatrix}zI_{n}&0\\ 0&wI_{m}\end{pmatrix}$ (4)
where $n+m=N$. This function is a rational inner function in two variables.
Conversely, the original proof of necessity in this theorem used a cone separation argument and Andô’s inequality from operator theory; see PickBook ; AM99 ; Agler1 . It was perhaps the insight of Cole-Wermer CW99 that it is enough to prove a sum of squares formula of the form
$|p(z,w)|^{2}-|\tilde{p}(z,w)|^{2}=(1-|z|^{2})\sum_{j=1}^{n}|E_{j}(z,w)|^{2}+(1-|w|^{2})\sum_{j=1}^{m}|F_{j}(z,w)|^{2}$ (5)
for $p\in\mathbb{C}[z,w]$ with no zeros in $\mathbb{D}^{2}$, and $E_{j},F_{j}\in\mathbb{C}[z,w]$. The moment problem approach mentioned above can be used to establish this formula independent of operator theoretic tools—this was done in Geronimo-Woerdeman GW and we discussed this connection further in gKBS . It turns out that formulas like (3) are interchangeable with sums of squares formulas as in (5). This equivalence is explained in detail in gKKummert . Kummert Kummert proved that all rational inner functions in two variables satisfy a formula like (3) using an approach motivated by control engineering, and thus gave an approach to much of this material that predates much of the mathematics literature. We gave a mathematical exposition of Kummert’s work in gKKummert .
It turns out to be quite interesting to develop and refine the formula (5) further. It will allow us to prove things about polynomials with no zeros on $\mathbb{D}^{2}$ and lead to additional new questions that we discuss in the next section.
For now, let us remark that currently there is no known interpolation theorem for $\mathbb{D}^{d}$ when $d>2$. An ‘interpolation theorem’ would mean that we reduce interpolation for bounded analytic functions to searching over a finite dimensional set. Theorem 0.3.1 does this even if the conditions given could be difficult to check. The following is not even known:
Problem 1
For $d>2$, given $z_{1},\dots,z_{N}\in\mathbb{D}^{d}$ and $f:\mathbb{D}^{d}\to\mathbb{D}$ analytic, does there exist a rational inner function $\phi$ in $d$ variables with $f(z_{j})=\phi(z_{j})$ for $j=1,\dots,N$?
The answer to this question is affirmative if $N\leq 3$ by work in Kosinski , gK3x3 .
0.4 Sums of squares
Using either a moment problem approach (GW , gKBS ) or Kummert’s approach (Kummert , gKKummert ), it is possible to prove the following refined version of (5).
Theorem 0.4.1
Let $\phi(z,w)=\frac{\tilde{p}(z,w)}{p(z,w)}$ be a rational inner function, where $p\in\mathbb{C}[z,w]$ has no zeros in $\mathbb{D}^{2}$, no factors in common with $\tilde{p}(z,w)$, and $\tilde{p}$ has bidegree $(n,m)$. Then, there exist polynomials $E_{1},\dots,E_{n}$ of bidegree at most $(n-1,m)$ and $F_{1},\dots,F_{m}$ of bidegree at most $(n,m-1)$ such that
$|p(z,w)|^{2}-|\tilde{p}(z,w)|^{2}=(1-|z|^{2})\sum_{j=1}^{n}|E_{j}(z,w)|^{2}+(1-|w|^{2})\sum_{j=1}^{m}|F_{j}(z,w)|^{2}.$
Equivalently, there exists a $(1+n+m)\times(1+n+m)$ unitary $U$ such that (3) holds.
The key thing is that the numbers of squares are precisely related to the bidegree of $\tilde{p}$. This precision is important in one of the main theorems of AMY on generalizing Loewner’s theorem on matrix monotone functions to several variables. It is also crucial to the following corollary, which gives a straightforward way to construct all polynomials with no zeros in $\mathbb{D}^{2}$.
Corollary 1
Suppose $p\in\mathbb{C}[z,w]$ has no zeros in $\mathbb{D}^{2}$ and bidegree $(n,m)$. Then, there exists an $(n+m)\times(n+m)$ contractive matrix $D$ such that
$p(z,w)=p(0,0)\det(I-D\Delta(z,w))$
with $\Delta$ as in (4).
If $p$ has no factors in common with $\tilde{p}$, the matrix $D$ can be given with additional structure— it can be chosen to be a rank one perturbation of a unitary; see Section 9 of gKextreme . In fact, polynomials with no zeros in $\mathbb{D}^{2}$ can be factored into $p=p_{1}p_{2}$ where $p_{1}$ has no factors in common with $\tilde{p}_{1}$ while $p_{2}$ is a constant multiple of $\tilde{p}_{2}$. In the latter case we can find a determinantal representation for $p_{2}$ where $D$ is in fact a unitary. Note that a direct proof of Corollary 1 is given in KummertStable , DrexelStable .
Returning to Theorem 0.4.1, it is worth pointing out that such sums of squares formulas are in general not unique. A trivial way they are not unique is that we could replace
$\begin{pmatrix}E_{1}\\ \vdots\\ E_{n}\end{pmatrix}\text{ and }\begin{pmatrix}F_{1}\\ \vdots\\ F_{m}\end{pmatrix}$
with unitary multiples. The sums of squares terms can be non-unique in a non-trivial way as well.
Example 2
For $p_{1}=3-z-w$ there exist distinct real numbers $A,B$ (one can solve for them) so that
$|p_{1}(z,w)|^{2}-|\tilde{p}_{1}(z,w)|^{2}$ $=(1-|z|^{2})|A+Bw|^{2}+(1-|w|^{2})|B+Az|^{2}$
$=(1-|z|^{2})|B+Aw|^{2}+(1-|w|^{2})|A+Bz|^{2}.$
Thus, this example has two inequivalent sums of squares formulas. $\diamond$
Understanding unique sums of squares decompositions or describing all such decompositions turns out to be surprisingly deep. It forces us to confront polynomials with boundary zeros. The following was proven in gKnoz (Theorem 1.15), gKintreg (Corollary 13.6).
Theorem 0.4.2
Let $p\in\mathbb{C}[z,w]$ have no zeros in $\mathbb{D}^{2}$, no factors in common with $\tilde{p}$, and $\deg\tilde{p}=(n,m)$. Consider the sums of squares terms $\mathcal{E}(z,w)=\sum_{j=1}^{n}|E_{j}(z,w)|^{2}$, $\mathcal{F}(z,w)=\sum_{j=1}^{m}|F_{j}(z,w)|^{2}$ appearing in Theorem 0.4.1. Then, the following are equivalent
  1. 1.
    the sums of squares terms $\mathcal{E},\mathcal{F}$ are unique (up to unitary transformation of the sums of squares)
  2. 2.
    the sums of squares terms possess the symmetries
    $\mathcal{E}(z,w)=|z^{n-1}w^{m}|^{2}\mathcal{E}(1/\bar{z},1/\bar{w})\quad\mathcal{F}(z,w)=|z^{n}w^{m-1}|^{2}\mathcal{F}(1/\bar{z},1/\bar{w})$
  3. 3.
[…]
  4. 4.
    $p$ and $\tilde{p}$ have $2nm$ common zeros on $\mathbb{T}^{2}$ counted using intersection multiplicities.
Item 2 above is related to the fact that every $p$ has two special sums of squares decompositions. Loosely speaking, one such decomposition has the property that $\mathcal{E}$ is maximal among decompositions and the other has $\mathcal{F}$ maximal. Item 2 then gives a condition that makes these two special decompositions coincide—and hence all decompositions coincide. See Theorem 1.3 of gKnoz .
The third and fourth items express the idea that $p$ has the maximal number of zeros on $\mathbb{T}^{2}$ without having a curve of zeros on $\mathbb{T}^{2}$. We called such polynomials saturated in gKextreme . The presence of the $L^{2}$ condition in item 3 is related to the construction of the sums of squares terms in gKnoz using bases for certain subspaces of $L^{2}(\frac{1}{|p|^{2}}d\sigma)$. Regarding item 4, Bézout’s theorem for $\mathbb{C}_{\infty}\times\mathbb{C}_{\infty}$ would say that $p$ and $\tilde{p}$ have $2nm$ common zeros globally, and item 4 is the condition that all of these zeros occur on $\mathbb{T}^{2}$. Again, common zeros are counted using intersection multiplicity as in elementary algebraic curve theory as in Fulton .
Example 3
The polynomial $p_{0}=2-z-w$ has unique sums of squares decomposition
$|p_{0}(z,w)|^{2}-|\tilde{p}_{0}(z,w)|^{2}=(1-|z|^{2})2|1-w|^{2}+(1-|w|^{2})2|1-z|^{2}.$
This can be checked using any of the above conditions 2-4 but symmetry is probably easiest to use. $\diamond$
It is probably not feasible to characterize all sums of squares decompositions, but it turns out it is possible to characterize the decompositions where the number of squares, $n$ for $\mathcal{E}$ and $m$ for $\mathcal{F}$, are minimal as in Theorem 0.4.1. The characterization is too technical to state here (see Theorem 9.3 of gKintreg ) but the most interesting aspect is that the minimal decompositions are in one-to-one correspondence with the invariant subspaces of a special pair of commuting contractive matrices.
Assume $p$ is given as in Theorem 0.4.1. Let
$\mathcal{G}=\left\{q\in\mathbb{C}[z,w]\cap L^{2}(\frac{1}{|p|^{2}}d\sigma):\deg q\leq(n-1,m-1)\right\}.$
Let $P$ denote orthogonal projection onto $\mathcal{G}$ in $L^{2}(\frac{1}{|p|^{2}}d\sigma)$. Consider the operators $T_{z},T_{w}:\mathcal{G}\to\mathcal{G}$ given by
$T_{z}q=P(zq)\qquad T_{w}q=P(wq).$
It is a non-trivial fact that $T_{z}$ and $T_{w}^{*}$ commute and the minimal decompositions of $p$ are in correspondence with the joint invariant subspaces of $(T_{z},T_{w}^{*})$. This result has important antecedents in work of Ball-Sadosky-Vinnikov BSV and Geronimo-Woerdeman GW .
Understanding sums of squares formulas for rational inner functions naturally leads to several questions. First, can we better understand the space
$\mathbb{C}[z,w]\cap L^{2}\left(\frac{1}{|p|^{2}}d\sigma\right)\text{ ?}$
Second, what happens in higher dimensions? The first question forces us to confront a local theory of stable polynomials (those with no zeros on $\mathbb{D}^{d}$) near a boundary zero. Along the way we learn much about the geometry of level sets of rational inner functions on $\mathbb{T}^{2}$. We explore this in the next section. The second question is inextricably linked to operator theory (the failure of von Neumann’s inequality in 3 or more variables). We explore this in the final section.
0.5 Boundary behavior
In this section we are interested in the behavior of rational functions $q/p$ near a boundary zero. A variety of behaviors could be studied including existence of non-tangential boundary values (or derivatives) as well as boundedness or integrability near/on the boundary.
Non-tangential behavior of general bounded analytic functions on polydisks has been studied a great deal in recent years. See Abate , AMYcara , AMYbook , MPcara , ATYboundary , Pascoe-cara , Tully-Doyle . However, there are some notable simplifications that occur for rational and rational inner functions. A main theorem of gKintreg , BKPS is the following.
Theorem 0.5.1
Suppose $p,q\in\mathbb{C}[z_{1},\dots,z_{d}]$, $p$ has no zeros on $\mathbb{D}^{d}$, and $f=q/p$ is bounded on $\mathbb{D}^{d}$. Then, $f$ has a non-tangential limit at every point of $\mathbb{T}^{d}$.
A non-tangential limit here just means that we approach a boundary point in such a way that the distance to the point is comparable to the distance to the boundary. A first step in proving this and related local problems is to switch to the setting of the poly-upper half plane
$\mathbb{H}^{d}=\{x=(x_{1},\dots,x_{d})\in\mathbb{C}^{d}:\Im x_{1},\dots,\Im x_{d}>0\}$
via conformal mapping where a boundary singularity is sent to the origin. For instance, the polynomial $p_{0}(z,w)=2-z-w$ converts to
$(1-ix)(1-iy)p_{0}\left(\frac{1+ix}{1-ix},\frac{1+iy}{1-iy}\right)=-2i(x+y-2ixy).$
Reflection in this setting becomes conjugation of coefficients:
$\bar{p}(x,y)=\overline{p(\bar{x},\bar{y})}.$
Thus, the favorite rational inner function $\frac{2zw-z-w}{2-z-w}$ becomes (after disregarding constants in front)
$\frac{x+y+2ixy}{x+y-2ixy}.$ (6)
The key local result needed for proving Theorem 0.5.1 is the following:
Theorem 0.5.2
Suppose $p\in\mathbb{C}[x_{1},\dots,x_{d}]$ has no zeros in $\mathbb{H}^{d}$ and $p(0)=0$. Write out the homogeneous expansion of $p(x)=\sum_{j=M}^{N}P_{j}(x)$ where $P_{j}$ is homogeneous of degree $j$ and $P_{M}\not\equiv 0$. Writing $P_{M+1}=A_{M+1}+iB_{M+1}$ for $A_{M+1},B_{M+1}\in\mathbb{R}[x_{1},\dots,x_{d}]$ we have the following.
  1. 1.
    $P_{M}$ has no zeros in $\mathbb{H}^{d}$ and there exists $\mu\in\mathbb{T}$ such that $\mu P_{M}\in\mathbb{R}[x_{1},\dots,x_{d}]$.
  2. 2.
    Assuming $\mu=1$ in part (a) and $p$ has no factors in common with $\bar{p}$, we have $B_{M+1}\not\equiv 0$ and $\frac{A_{M}}{B_{M+1}}$ maps $\mathbb{H}^{d}$ into $\mathbb{H}$.
This is stated as the “Homogeneous Expansion” theorem in BKPS where the first part is attributed to ABG . The next step is proving that a rational function with given homogeneous expansions
$\frac{\sum_{j=K}^{L}Q_{j}(x)}{\sum_{j=M}^{N}P_{j}(x)}$ (7)
has a non-tangential limit at 00 if and only if $K=M$ and $Q_{K}$ is a multiple of $P_{M}$. This property is evident in the example (6) where $M=1$ and $P_{1}=x+y$. Finally, one proves that this property holds for bounded rational functions.
Next, we consider boundedness as well as boundary integrability near a boundary point. Boundedness is interesting because of the natural question of simply characterizing the bounded rational functions. Square integrability is of interest because of the previous section where certain sums of squares formulas depend on spaces of square integrable rational functions. Based on the above, one might hope that boundedness near a boundary point (meaning boundedness on $\mathbb{H}^{d}$ intersected with a neighborhood of 00) would simply depend on the condition that $Q_{M}=cP_{M}$ as in (7). Unfortunately, this is only true in the simplest examples like (6).
Example 4
The polynomial
$p_{2}(x,y)=x+y-ix^{2}-ixy-2x^{3}-6x^{2}y+4ix^{3}y$
has no zeros in $\mathbb{H}^{2}$ while
$\frac{x+y}{p_{2}(x,y)}$
is unbounded near $(0,0)$. Indeed, we can show it is unbounded on $\mathbb{R}^{2}\setminus\{(0,0)\}$ near $(0,0)$ and then it will be unbounded on $\mathbb{H}^{2}$ near 00 by continuity. Setting $y=-x-4x^{3}$ we have
$\frac{-4x^{3}}{24x^{5}-16ix^{6}}=\frac{1}{-6x^{2}+4ix^{3}}$
which goes to $\infty$ as $x\to 0$. This is Example 5.6 of BKPS . $\diamond$
To understand what is going on we look at the Puiseux series for the branches of $p(x,y)$ with no zeros in $\mathbb{H}^{2}$ but $p(0,0)=0$. Recall that by the Newton-Puiseux theorem, the zero set of $p$ near $(0,0)$ can be parametrized (injectively) by maps of the form
$t\mapsto(t^{n},\phi(t))$
where $n$ is a positive integer and $\phi$ is analytic in a neighborhood of $0\in\mathbb{C}$ with $\phi(0)=0$. This leads to a factorization of $p$ with factors of the form
$\prod_{j=1}^{n}(y-\phi(\mu^{j}x^{1/n}))$
where $\mu=\exp(2\pi i/n)$. (See simonbasic , Theorems 3.5.1, 3.5.2.) In the current case, we have the property that whenever $t^{n}\in\mathbb{H}$ then $\phi(t)\notin\mathbb{H}$ (i.e. $\phi(t)\in-\overline{\mathbb{H}}$ the closed lower half plane). It turns out that we can characterize this. It helps to replace $\phi$ with its negative, so that our characterization has positive parameters.
Theorem 0.5.3
Suppose $\phi(t)$ is analytic in a neighborhood of $0\in\mathbb{C}$ and $\phi(0)=0$. Let $n$ be a positive integer and suppose that
$t\mapsto(t^{n},-\phi(t))$
is an injection into $\mathbb{C}^{2}\setminus\mathbb{H}^{2}$. Then, there are two cases:
  1. 1.
    (real stable type) $n=1$, $\phi$ has real coefficients and $\phi^{\prime}(0)>0$.
  2. 2.
    (pure stable type) There exist a positive integer $L$, $q\in\mathbb{R}[t]$ with $\deg q<2L$, $q(0)=0$, $q^{\prime}(0)>0$, and $\psi(t)$ analytic in a neighborhood of 00 with $\Im(\psi(0))>0$ such that
    $\phi(t)=q(t^{n})+t^{2Ln}\psi(t).$
In the first case, $p(x,y)$ will have an analytic factor $y+\phi(x)$ with $\phi^{\prime}(0)>0$ where $\phi$ has real coefficients. This implies $p$ has a smooth/analytic curve of zeros in the distinguished boundary $\mathbb{R}^{2}$. In the second case, $p(x,y)$ will have a factor of the form
$\prod_{j=1}^{n}(y+q(x)+x^{2L}\psi(\mu^{j}x^{1/n})).$
We will prove the theorem here because its original statement (as Lemma C.3 of gKintreg ) neglected the real coefficient case (as it was not relevant in that context).
Proof
Let us write out the initial term of $\phi(t)=at^{r}+\dots$ with $a\neq 0$. We first show that $r$ equals $n$ and $a>0$. Writing $t=|t|e^{i\theta}$, we have that for $n\theta\in(0,\pi)+2\pi\mathbb{Z}$, $\phi(|t|e^{i\theta})\in\overline{\mathbb{H}}$. Writing $a=|a|e^{i\alpha}$ we have
$\lim_{|t|\searrow 0}\frac{\phi(|t|e^{i\theta})}{|t|^{r}}=|a|e^{i\alpha}e^{ir\theta}$
and therefore $n\theta\in(0,\pi)+2\pi\mathbb{Z}$ implies $\alpha+r\theta\in[0,\pi]+2\pi\mathbb{Z}$.
This means that intervals of the form $(\alpha+(2j)\frac{\pi r}{n},\alpha+(2j+1)\frac{\pi r}{n})$ for $j\in\mathbb{Z}$ must be contained in $[0,\pi]+2\pi\mathbb{Z}$. Since the width of the open intervals is $r\pi/n$ we must have $r\pi/n\leq\pi$ so that $r\leq n$. The contrapositive says if $\alpha+r\theta\in(\pi,2\pi)+2\pi\mathbb{Z}$ then $n\theta\in[\pi,2\pi]+2\pi\mathbb{Z}$ and a similar argument shows $n\leq r$. In order to have $(\alpha,\alpha+\pi)\subset[0,\pi]+2\pi\mathbb{Z}$ we must have $\alpha$ an even multiple of $\pi$. Therefore, $a=|a|e^{\alpha}>0$.
Next we note that by continuity we must have that whenever $t^{n}\in\mathbb{R}$ then $\phi(t)\in\overline{\mathbb{H}}$. Writing $\phi(t)=\sum_{j=n}^{\infty}a_{j}t^{j}$, we define
$\phi_{0}(t)=\sum_{j=1}^{\infty}\Re(a_{jn})t^{jn}=at^{n}+\Re(a_{2n})t^{2n}+\cdots$
which just extracts the real coefficients of powers $t^{jn}$. If $\phi(t)=\phi_{0}(t)$, then $t\mapsto(t^{n},\phi(t))$ is not injective unless $n=1$. Thus, in this case $\phi$ has the first form of the theorem.
So, suppose $\phi(t)\neq\phi_{0}(t)$. Since $t^{n}\in\mathbb{R}$ implies $\phi_{0}(t)\in\mathbb{R}$, the function
$\tilde{\phi}(t)=\phi(t)-\phi_{0}(t)$
has the property that $t^{n}\in\mathbb{R}$ implies $\tilde{\phi}(t)\in\overline{\mathbb{H}}$. Let us write $\tilde{\phi}(t)=\sum_{j=m}^{\infty}b_{j}t^{j}=bt^{m}+\cdots$, where $b=b_{m}=|b|e^{i\beta}\neq 0$ and $m\geq 2$. Writing $t=|t|e^{i\theta}$ and taking a limit as before we see that $n\theta\in\pi\mathbb{Z}$ implies $\beta+m\theta\in[0,\pi]+2\pi\mathbb{Z}$. In other words, after dividing by $\pi$
$\beta/\pi+\frac{m}{n}\mathbb{Z}\subset[0,1]+2\mathbb{Z}.$
By an elementary lemma in gKint (Lemma 3.4), this implies $m/n$ is an integer and if $m/n$ is odd then $\beta/\pi\in\{0,1\}$. The latter condition would mean the first term of $\tilde{\phi}(t)$ has the form $bt^{(2L+1)n}$ with $b\in\mathbb{R}$. But $\phi_{0}$ removed all such terms, so we must have $m=2Ln$ for some $L$. In this case, $\tilde{\phi}(t)$ begins $bt^{2Ln}$ and $b$ must be purely imaginary (and nonzero) since $\phi_{0}$ removed the real parts of such terms. Since $\beta\in[0,\pi]$, we must have $\beta=\pi/2$; namely $b$ is a positive multiple of $i$. Thus,
$\phi(t)=\phi_{0}(t)+\tilde{\phi}(t)$
which we can write as $q(t^{n})+t^{2Ln}\psi(t)$ with $q$ and $\psi$ as in the theorem statement. ∎
Example 5
The polynomial $p=x+y-2ixy$ factors as
$(1-2ix)\left(y+\frac{x}{1-2ix}\right)=(1-2ix)(y+x+2ix^{2}+\text{ higher order})$
and therefore $\phi(x)=\frac{x}{1-2ix}=x+x^{2}\frac{2i}{1-2ix}$ is an example of the theorem with $q(x)=x$, $L=1$, $n=1$, $\psi(x)=\frac{2i}{1-2ix}$. More complicated examples are presented in BKPS $\diamond$
If $p(x,y)$ has no zeros in $\mathbb{H}^{2}$ and no factors in common with $\bar{p}$, then all of $p$’s branches are of the second type called pure stable type. This is because branches of the first type (real stable type) have infinitely many common zeros on $\mathbb{R}^{2}$ and this would imply $p$ and $\bar{p}$ have a common factor by Bézout’s theorem. We obtain the following complete local description of such $p$. We use the notation $\mathbb{C}\{x,y\}$ to denote the set of power series in $x,y$ that converge in a neighborhood of the origin.
Theorem 0.5.4 (Theorem 2.16 in BKPS )
Suppose $p\in\mathbb{C}[x,y]$ has no zeros in $\mathbb{H}^{2}$, no factors in common with $\bar{p}$, and $p(0,0)=0$. Then, there exist $u\in\mathbb{C}\{x,y\}$ with $u(0,0)\neq 0$, positive integers $2L_{1},\dots,2L_{k}$, $M_{1},\dots,M_{k}$, polynomials $q_{1},\dots,q_{k}\in\mathbb{R}[x]$ where
  • $\deg q_{j}<2L_{j}$
  • $q_{j}(0)=0$, $q_{j}^{\prime}(0)>0$
and $\psi_{1},\dots,\psi_{k}\in\mathbb{C}\{x\}$ with $\Im\psi_{j}(0)>0$ such that
$p(x,y)=u(x,y)\prod_{j=1}^{k}\prod_{m=1}^{M_{j}}(y+q_{j}(x)+x^{2L_{j}}\psi_{j}(\mu_{j}^{m}x^{1/M_{j}}))$
where $\mu_{j}=\exp(2\pi i/M_{j})$.
The data $2L_{1},\dots,2L_{k},M_{1},\dots,M_{k},q_{1}(x),\dots,q_{k}(x)$ is the most important local information about $p$. In fact, if we define
$[p](x,y)=\prod_{j=1}(y+q_{j}(x)+ix^{2L_{j}})^{M_{j}}$ (8)
then Theorem 1.2 of BKPS says that
$\frac{[p](x,y)}{p(x,y)}$
is bounded above and below near 00 in $\mathbb{H}^{2}$. Thus, for many problems we can replace $p$ with the simpler polynomial $[p]$. At the same time, Theorem 0.5.4 makes it possible to construct polynomials with no zeros in $\mathbb{H}^{2}$ or $\mathbb{D}^{2}$ with non-trivial Puiseux structure at a boundary point—something that was previously not known. The construction amounts to writing out a proposed local decomposition such as
$P(x,y)=(y+x+ix^{2}+x^{5/2})(y+x+ix^{2}-x^{5/2})=(y+x+ix^{2})-x^{5}$
which one can show has no zeros in the $\mathbb{H}^{2}\cap\mathbb{D}_{\epsilon}^{2}$. Then, we can find a pair of conformal maps $\phi_{1}(x),\phi_{2}(y)$ that map $\mathbb{H}$ to a disk in $\mathbb{H}$ that is tangent to $\mathbb{R}$ at 00 so that $Q=P(\phi_{1}(x),\phi_{2}(y))$ has no zeros in $\mathbb{H}^{2}$ yet we do not change the fact that a local factorization of $Q$ requires Puiseux series. We can clear the denominators of $Q$ to get a polynomial non-vanishing on $\mathbb{H}^{2}$ with non-trivial Puiseux structure at $(0,0)$. This process blurs the original data associated with $P$, so we pose it as a problem later (Problem 2) to show that we can construct global stable polynomials with prescribed local data.
Theorem 0.5.4 also led to the following conjecture and now theorem characterizing rational functions bounded near a boundary point.
Theorem 0.5.5 (Theorem 5.2 of BKPS , Theorem 3 of kollar )
Assume $p$ as in Theorem 0.5.4. Given $q\in\mathbb{C}[x,y]$, the rational function $q/p$ is bounded in a neighborhood of $(0,0)$ intersected with $\mathbb{H}^{2}$ if and only if $q$ belongs to the product ideal
$\prod_{j=1}^{k}(y+q_{j}(x),x^{2L_{j}})^{M_{j}}$
in $\mathbb{C}\{x,y\}$.
A different proof of this theorem was given in gKint . For example, for $p=x+y-2ixy$, the numerators for locally bounded rational functions belong to the ideal generated by $(y+x,x^{2})$.
Important prior work related to Theorems 0.5.4 and 0.5.5 was done in the papers BPS1 , BPS2 where the local geometry of the zero set of $p$ was connected to the local geometry of the level sets of rational inner functions now written in the upper half plane setting as $\phi(x,y)=\bar{p}(x,y)/p(x,y)$
$\{(x,y)\in\mathbb{R}^{2}:\phi(z,w)=\alpha\}\qquad\alpha\in\mathbb{T}.$ (9)
The data $2L_{1},\dots,2L_{k}$ from Theorem 0.5.4 can be used to measure the rate that the zero set of $p$ approaches $\mathbb{R}^{2}$ near $(0,0)$. Specifically, $K_{0}=\max\{2L_{1},\dots,2L_{k}\}$ is called the local contact order of $p$ at $(0,0)$ because for real $x$ near 00 and $\epsilon>0$ sufficiently small
$\inf\{|\Im y|:p(x,y)=0,|y|<\epsilon\}\approx|x|^{K_{0}}.$
Contact order is conformally invariant and in the setting of $\mathbb{D}^{2}$ and $\mathbb{T}^{2}$ one can define a global contact order $K$ by taking the maximum of local contact orders at all boundary singularities. A main result of Bickel-Pascoe-Sola BPS1 ,BPS2 is the following characterization of the integrability of derivatives of rational inner functions.
Theorem 0.5.6
Let $\phi(z,w)=\tilde{p}(z,w)/p(z,w)$ be a rational inner function on $\mathbb{D}^{2}$ with global contact order $K$. Then, for $1\leq\mathfrak{p}<\infty$
$\frac{\partial\phi}{\partial z}\in L^{\mathfrak{p}}(\mathbb{T}^{2})\text{ if and only if }\frac{\partial\phi}{\partial w}\in L^{\mathfrak{p}}(\mathbb{T}^{2})\text{ if and only if }\mathfrak{p}<1+\frac{1}{K}.$
Crucial to the approach of BPS1 , BPS2 was showing that the branches of the level sets (9) occur in bundles called horn regions. Figure 1 shows various level sets of a rational inner function. Note how they occur in two horn-shaped regions or bundles. These bundles are directly connected to the polynomials $q_{1}(x),\dots,q_{k}(x)$ in Theorem 0.5.4. Necessarily, for a rational inner function $\phi$, we have $\frac{\partial\phi}{\partial z}$ belonging to $L^{1}(\mathbb{T}^{2})$. This fact even extends to several variables. In two variables, the paper Pascoe-cara shows how to construct polynomials with given global contact order so that conditions in Theorem 0.5.6 can actually occur.
Figure 1: Various level sets for a rational inner function
Finally, in the preprint gKint we gave a complete characterization of when a rational function $Q/p$ (where $p$ has no zero in $\mathbb{H}^{2}$ and no factors in common with $\bar{p}$) is locally in $L^{\mathfrak{p}}$ in $\mathbb{R}^{2}$ near the origin. The main result involves checking the behavior of $Q$ along curves related to the polynomials $q_{1}(x),\dots,q_{k}(x)$ from Theorem 0.5.4. A few of the easiest to state offshoots are below. In the next theorem, we consider the space of $Q/p$ that are locally square integrable:
[…]
Notice that both $p$ and $\bar{p}$ belong to $\mathcal{I}_{p}$ so it makes sense to quotient $\mathcal{I}_{p}$ by the ideal generated by $p$, $\bar{p}$; namely $(p,\bar{p})$.
Theorem 0.5.7
Assume $p$ satisfies the setup and conclusion of Theorem 0.5.4. Then,
$\dim\mathcal{I}_{p}/(p,\bar{p})=\frac{1}{2}\dim\mathbb{C}\{x,y\}/(p,\bar{p}).$
Another easy to state result related to Theorem 0.5.6 is the following.
Theorem 0.5.8 (Theorem 1.9 of gKint )
Suppose $f:\mathbb{D}^{2}\to\mathbb{C}$ is bounded and rational. Then, $\frac{\partial f}{\partial z},\frac{\partial f}{\partial w}$ belong to $L^{1}(\mathbb{T}^{2})$.
Before we leave two variables, it worth highlighting a few other areas where rational inner functions have appeared. Recently, several authors have begun studying the Clark measures associated to rational inner functions Doubtsov , Clark1 , Clark2 . This work involves a detailed study both of the sums of squares decompositions of the previous section as well as the local geometry from the present section. In addition, the dynamics of rational inner mappings (pairs of rational inner functions) have been studied in dyn1 , dyn2 . One interesting by-product of dyn2 is that by composing rational inner functions with singularities one can produce polynomials with high contact order. This leads to a question.
Problem 2
Given the data $2L_{1},\dots,2L_{k},M_{1},\dots,M_{k},q_{1}(x),\dots,q_{k}(x)$ from Theorem 0.5.4 can one always construct a polynomial $p\in\mathbb{C}[x,y]$ with no zeros in $\mathbb{D}^{2}$ associated to this data as in Theorem 0.5.4?
0.6 Higher dimensions
In this final section we look at what is known about rational inner functions in three or more variables. As indicated in Section 0.3, if one could understand interpolation for rational inner functions, it ought to be possible to understand interpolation more broadly but if the answer to Problem 1 is negative, then a different approach may be necessary.
It turns out that the natural generalizations of Agler’s Pick interpolation (Theorem 0.3.1) and the sums of squares theorem (Theorem 0.4.1) do not hold for three or more variables, and this is well-known to be connected to operator theoretic obstructions. An analytic function $f:\mathbb{D}^{d}\to\mathrm{d}$ satisfies a formula
$1-f(z)\overline{f(w)}=\sum_{j=1}^{d}(1-z_{j}\bar{w}_{j})K_{j}(z,w)$ (10)
where the $K_{j}$ are positive semi-definite kernels if and only if $f$ satisfies an operator inequality: for every $d$-tuple of commuting strictly contractive operators $T=(T_{1},\dots,T_{d})$ on a Hilbert space, we have $\|f(T)\|\leq 1$. For $d=1$, this inequality follows automatically—this is von Neumann’s inequality vN —and for $d=2$ it follows automatically by Andô’s inequality Ando . For $d>2$, it does not follow automatically because there are counterexamples (first discovered by Varopoulos varo ). For this reason it makes sense to study the class of functions for which the operator inequality holds.
Definition 2
We say analytic $f:\mathbb{D}^{d}\to\mathbb{D}$ belongs to the Schur-Agler class (or Agler class for short) if for every $d$-tuple of commuting strictly contractive operators $T=(T_{1},\dots,T_{d})$ we have $\|f(T)\|\leq 1$.
In particular, we can study rational inner functions in the Schur-Agler class.
Theorem 0.6.1
A rational inner function $\phi(z)=\tilde{p}(z)/p(z)$ on $\mathbb{D}^{d}$ belongs to the Schur-Agler class if and only if we have a formula
$|p(z)|^{2}-|\tilde{p}(z)|^{2}=\sum_{j=1}^{d}(1-|z_{j}|^{2})\mathcal{E}_{j}(z)$
where each $\mathcal{E}_{j}(z)$ is a sum of squared moduli of polynomials.
The concept of this theorem is essentially due to Cole-Wermer CW99 , while it is explicitly stated in gKratag with explicit degree bounds for the terms in the sums of squares terms. Certain low-degree or regular rational inner functions automatically belong to the Agler class.
  • If $p(z_{1},z_{2},z_{3})$ with no zeros in $\mathbb{D}^{3}$ has multidegree $(n,1,1)$ then $\tilde{p}/p$ is in the Agler class; this is proven in gKtridisk with some elaborations in BK .
  • If $p(z_{1},z_{2},z_{3})$ has multidegree $(n,m,1)$ and no zeros in $\mathbb{D}^{3}$, then there is a monomial $z^{\alpha}$ such that $z^{\alpha}\tilde{p}/p$ belongs to the Agler class. See gKtridisk . (Note the theorem in gKtridisk states $p$ has no zeros in $\overline{\mathbb{D}}^{3}$ but this is not essential if we use a theorem of Scheiderer scheid on non-negative trigonometric polynomials.)
  • In drexelcon , it is proven that if $p\in\mathbb{C}[z_{1},\dots,z_{d}]$ has no zeros on $\overline{\mathbb{D}}^{d}$ then there is a monomial $z^{\alpha}$ such that $z^{\alpha}\tilde{p}/p$ belongs to the Agler class.
  • In the paper gKsym we gave a concrete checkable condition for when polynomials $p(z_{1},\dots,z_{d})$ with no zeros in $\mathbb{D}^{d}$ which are multiaffine (meaning they have degree one in every variable) and are symmetric give rise to $\phi=\tilde{p}/p$ belonging to the Agler class. The following is not known as far as we know.
Problem 3
Suppose $p(z_{1},\dots,z_{d})$ has no zeros in $\mathbb{D}^{d}$, is multiaffine and symmetric. Does $\phi=\tilde{p}/p$ automatically belong to the Agler class?
The answer is probably no, but if true it would represent a novel strengthening of the Grace-Walsh-Szegő theorem.
One can define an Agler norm for analytic $f:\mathbb{D}^{d}\to\mathbb{C}$ via
$\|f\|_{A}=\sup_{T}\|f(T)\|$
where the supremum is taken over commuting strictly contractive $d$-tuples. We could have $\|f\|_{A}=\infty$. A phenomenon noted in drexelschwarz is that there exists a rational inner function $\phi$ in three variables with $\|\phi\|_{A}>1$ and there exists a monomial $z^{\alpha}$ such that $\|z^{\alpha}\phi\|_{A}=1$.
A basic open question suggested by V. Vinnikov is the following:
Problem 4
Does every rational inner function have finite Agler norm?
By absolute convergence properties, if $p$ has no zeros on $\overline{\mathbb{D}}^{d}$ then necessarily $\tilde{p}/p$ has finite Agler norm so the question is about rational inner functions with boundary singularities. This leads to further motivation for studying boundary singularities of rational functions.
The paper BPS3 presents a detailed study of the boundary level sets of rational inner functions in higher dimensions. One of the key results demonstrates that integrability of the derivative of a rational inner function depends on the rate that the zero set of $p$ approaches the boundary. This paper also presents a variety of interesting examples where the zero set of $p$ intersected with $\mathbb{T}^{3}$ has either an isolated zero or a curve of singularities. There seems to be no obvious connection between coarse boundary nature and derivative integrability. The paper Bergqvist has some further results on integrability in higher dimensions. A basic question coming out of BPS3 is the following.
Problem 5
The exact derivative integrability cut-offs are known for rational inner functions in two variables by Theorem 0.5.6. What are the possible integrability cut-offs in three or more variables? Specifically, given a rational inner function $\phi$ on $\mathbb{D}^{d}$ for $d>2$, what are all of the possible values of
$\sup\left\{\mathfrak{p}>0:\frac{\partial\phi}{\partial z_{1}}\in L^{\mathfrak{p}}(\mathbb{T}^{d})\right\}\text{ ?}$
Finally, the preprint BKPS2 studies the problem of characterizing $q/p$ in three variables that are bounded near a boundary singularity in the simplest case, namely when the zero set of $p$ is smooth at the boundary singularity. There remain many open questions in three or more variables.
0.7 Acknowledgements
I would like to thank Alan Sola for offering comments on a draft of this paper (on extremely short notice!). Also, thanks to all the collaborators and friends who have contributed to the interesting (to me) topic of rational inner functions. Finally, thank you to the anonymous referee for numerous helpful suggestions. An additional thanks goes to Rongwei Yang for pointing out a typo in the proof of Prop 2.